### 26 Jan 2024 v1

@@ 2.3 Identification of bars
Non-axisymmetric structures are identified by Fourier decomposing the face-on stellar surface density (e.g., Athanassoula & Misiriotis 2002; Zana et al. 2018a; Rosas-Guevara et al. 2020). We focus on $A_{2}(R)$, which is defined by the ratio between the second and zero terms of the Fourier expansion and its phase $\Phi(R)$ (see equations 1 and 2 in RG22). Both quantities $A_{2}(R)$ and $\Phi(R)$, have been used to characterise a bar structure, where the bar strength is defined as the value of the peak of $A_{2}(R)$, $A_{\rm 2,max}$. The phase should be constant inside the bar. We define that the phase is constant by calculating the standard deviation ($\sigma$) of $\Phi(R)$, including each time a new cylinder shell, and imposing $\sigma\leq 0.1$. Then, we define the bar extent ($r_{\rm bar}$) as the maximum radius where $\sigma\leq 0.1$, and the A2 profile first dips at $0.15$ or the minimum value of $A_{2}(R)$. As large values of $A_{2}(R)$ could also be due to transient events, such as mergers and interactions, we conservatively assume the bar is a long-lasting feature if
@@
    The maximum of $A_{2}$, $A_{\rm 2,max}$, is greater than $0.2$,
@@
    The estimated age of the bar is larger than the time between the analysed output and 2 previous simulation outputs ($0.33$ Gyrs at $z=0$ and $0.17$ Gyrs at $z=4$).
At lower redshifts, this criterion may exclude recently produced bars but remains consistent with the bar selection criteria used by RG20, when the presence of a bar structure was observed in earlier outputs.
Besides the criterion to identify the bar structure, it is required to determine the formation time of a bar, $t_{\rm bar}$. To this end, we retrace the evolution of $A_{\rm 2,max}$ up to the lookback time when $A_{\rm 2,max}\leq 0.2$ and $r_{\rm bar}(t_{\rm bar})\leq r_{\rm min}$ for more than two snapshots. In addition, the relative difference between the $A_{\rm 2,max}$ at a given snapshot and the $A_{\rm 2,max}$ at the two prior snapshots must not exceed $0.45$ within this time frame. This provides additional assurance that the bars we detect are stable structures.
@@
The catalogue of RG22 consists of 260 disc galaxies at, $z=1$ of which $125$ (0.48±0.03plus-or-minus0.480.030.48$\pm$ 0.030.48 ± 0.03333These are the binomials errors on the bar fractions which are calculated using the number of galaxies and bars as $\sigma=(f_{\rm bars}(1-f_{\rm bars})/n_{\rm discs})^{0.5}$ ) have a bar. In this work, we ignore the galaxies that merged with a larger galaxy. This condition reduced our sample to $204$ disc galaxies at $z=1$ and with $93$ barred galaxies ($0.45\pm 0.03$). The details of the total sample can be found in Table 1. We follow the evolution of these galaxies from $z=1$ to $z=0$. From herein, we will tag the $z=0$ sample as descendant galaxies.
@@ 2.4 Descendant galaxy samples
In order to ascertain the primary mechanisms influencing the disc galaxies at $z=1$, we employ the Sublink merger tree technique (Rodriguez-Gomez et al. 2015) to track their evolution over time. The parent disc sample was partitioned into three distinct subsamples based on the final morphology of their descendant galaxies at $z=0$. Our analysis specifically centres on the stellar mass fraction of the cold disc component at $z=0$ ($(D/T)_{\rm thin,z0}$), as our initial dataset mostly consists of galaxies with a prominent cold disc component. The division was made in the following manner:
  • disc descendants refer to those galaxies at $z=1$ and their descendant galaxies at $z=0$ which have a massive cold disc component, defined as $(D/T)_{\rm thin,z0}\geq 0.5$ and ($(D/T)_{\rm thin,z1}\geq 0.5$).
  • lenticular descendants, defined as those galaxies at $z=1$ have a massive cold disc component $(D/T)_{\rm thin,z1}\geq 0.5$. These galaxies have descendants at $z=0$ that possess an intermediate cold disc component, specifically falling within the range of $0.50.3\geq(D/T)_{\rm thin,z0}<$ 0.5.
  • spheroidal descendants, defined as those galaxies at $z=1$ have a massive cold disc component $(D/T)_{\rm thin,z1}\geq 0.5$, however, their descendants at $z=0$ have seen a reduction in their cold disc component, with values smaller than $(D/T)_{\rm thin,z0}=0.3$.
The disc descendants comprise $104$ galaxies, representing $51$ per cent of the total parent sample, of which 36 galaxies are barred at $z=1$. Among these, there are $47$ galaxies classified as lenticular descendants ($23$ per cent of the parent samples) with $26$ barred galaxies at $z=1$. Finally, we get $53$ galaxies classified as spheroidal descendants, corresponding to $26$ per cent of the parent sample. The details can be seen in Table 1. In the evolution of the bar structures, we only focus on disc and lenticular galaxies.
@@ 2.5 Quantifying the effects of environment
The main aim of this study is to conduct a quantitative analysis of the impact of environmental factors on the development and progression of bars. Previous studies have provided evidence that mergers could facilitate the formation of a bar structure (e.g. Elmegreen, Elmegreen, & Bellin 1990; Łokas 2018). On the other hand, alternative studies have demonstrated that mergers can lead to the disruption of pre-existing bars (e.g. Zana et al. 2018b). Tidal interactions, such as those resulting from galaxy-galaxy interactions and flybys, are an additional external mechanism that can contribute to the formation of bars (e.g. Elmegreen, Elmegreen, & Bellin 1990; Lang, Holley-Bockelmann, & Sinha 2014; Łokas 2018; Peschken & Lokas 2019; Izquierdo-Villalba et al. 2022). This work aims to quantitatively assess three environmental factors, including the number of satellites as a function of time, that may potentially influence the evolution and production of bars. By identifying these external mechanisms, we seek to gain a deeper understanding of their impact on bars.
@@ 2.5.1 Merger histories and their influence
Major mergers are characterised by a stellar mass ratio of the secondary galaxy to the primary galaxy, denoted as $\mu$, that surpasses $0.25$, whereas minor mergers are defined by $\mu$ values ranging from $0.01$ to $0.25$  [^4] The lower limit of the mass ratio for minor mergers ensures that both galaxies are properly followed. Since our galaxy sample have stellar masses $>10^{10}\rm M_{\sun}$, the secondary galaxy will have $M_{*}>10^{8}\rm M_{\sun}$ which correspond to above $1000$ initial mass of the gas cells.. In order to quantify the parameter $\mu$, the maximum stellar masses of the galaxies in preceding snapshots are used, namely, those snapshots in which both galaxies are recognised as distinct structures by the SUBLINK algorithm.
According to McAlpine et al. (2020), a merging influence temporal zone can be defined as the period during which galaxies are expected to undergo a close merger within two dynamical timescales of the halo, relative to the desired lookback time (time of bar formation, time of bar dissolution). In this particular scenario, our focus lies on the temporal interval during which the bar undergoes formation or dissolution. The time period influenced by the merging process can be parameterized as
@@
where $t_{\rm obs}$ is the lookback time of formation/dissolution of the bar, $t_{\rm m}$, lookback time of the merger, and the dynamical time of the halo is defined as the free-fall time of a dark matter halo, i.e.,
$t_{\rm dyn,obs}\equiv\left(\frac{3\pi}{32G(200\rho_{\rm crit})}\right)^{1/2}$ (2)
This dynamical time of the halo depends on redshift. As a point of reference, the dynamical time of a halo for tree redshifts is approximately $0.9$ Gyr at $z=1$, $1.2$ Gyr at $z=0.5$, and $1.6$ Gyr at $z=0$. Positive values of $\eta_{\rm dyn,obs}$ indicate that the closest merger has already occurred, whereas negative values indicate that it will occur in the future. Absolute values of $|\eta_{\rm dyn,obs}|<1$ indicate that the merger takes place in one dynamical time. Therefore, we define the merger influence region for galaxies where the nearest merger occurs as $|\eta_{\rm dyn,obs}|<2$ whereas values higher than $|\eta_{\rm dyn,obs}|>2$ are not influenced by a merger. Here we will focus mainly on major mergers, as they have the greatest potential for significant impact. However, we also calculated this for minor mergers.
@@ 2.5.2 Close massive companions
In a similar vein, to measure galaxies with the likelihood of encountering a flyby event (not necessarily leading to a merger event), it is necessary to compute the proximity of the nearest large companion, where the ratio of their stellar masses exceeds $1/4$. Galaxies that may experience an influence due to a nearby encounter will possess nearby massive partners situated within a maximum distance of $d_{\rm cp}<100$ kpcs (Łokas 2021). Conversely, galaxies that are usually isolated are characterised by the closest massive companion being located at a distance greater than $d_{\rm cp}>500$ kpc.
@@ 3 Morphological evolution of $z=1$ disc galaxies
Figure 1: The stellar mass at of their descendants at $z=0$ as a function of the stellar mass of disc galaxies at $z=1$. Different markers and colours represent different morphology of the descendant galaxies, as the legend indicates. The black solid line represents one-to-one relation. Panels along the margins show the distributions of stellar mass at $z=0$ and $z=1$. Dotted lines represent the median of each distribution and the black dotted line is the median stellar mass of the total sample. There is a mild evolution in stellar mass in the last $8$ Gyrs in the parent disc sample.
Before analysing the evolution of the bar structures, we will be focusing on the morphological evolution of our parent sample, as defined in section 2.2.
@@
We investigate the evolution of our parent disc galaxies at $z=1$ by employing the tree merger algorithm outlined in section 2.1. The primary focus of the analysis is on the galaxy descendants of these disc galaxies at $z=0$.
@@
Fig 1 displays the scatter plot of the stellar mass of the parent disc sample at $z=1$ versus the stellar mass of their descendants at $z=0$. We can appreciate from the figure that, overall, there was an increase in stellar mass, although this growth is modest, with a median stellar mass of $10^{10.33}\rm M_{\sun}$ at $z=1$ rising to $10^{10.60}\rm M_{\sun}$ at $z=0$. We also note that some outliers, especially from spheroidal descendant galaxies, deviate from the overall distribution. The stellar mass of these galaxies decreases due to the entrance into a densely populated region, probably experiencing tidal stripping, as we will see in the next section. This is confirmed when we look at the stellar mass distributions (sided small panels) of the descendant galaxies split by shape, where the median of the stellar mass at $z=0$ is smaller in spheroidal compared to those of disc and lenticular descendants whereas the parent sample at $z=1$ shows that the progenitors of the lenticular and spheroidals were more massive.
@@
It is important to note that the morphological evolution of our parent sample exhibits significant divergence. Fig 2 displays the distribution in the disc stellar mass fraction, denoted as $D/T$, represented by the grey colour. The median $D/T$ decreases from $0.88$ at redshift $z=1$ to $0.75$ at redshift $z=0$. The dominant contribution is attributed to the thin disc component, which exhibits a median of $(D/T)_{\rm thin,z1}=0.67$ at $z=1$. Nevertheless, when considering the case where $z=0$, it can be shown that the distribution of $(D/T)_{\rm thin,z0}$ in the cold disc component is wider, spanning from $0.0$ to values higher than $0.75$. The median value of $(D/T)_{\rm thin,z0}$, is found to be $0.50$. This wider distribution of the cold component is also evident in the overall distribution of the discs, including all components. It should be noted that there is a lack of substantial evolutionary changes observed in the thick disc component, as indicated by a median value of $(D/T)_{\rm thin,z0,z1}\approx 0.2$ for both redshifts.
@@
In order to examine the changes in morphology within our parent sample, we have chosen three galaxy descendant samples as outlined in section 2.4 and Table 1. These samples have been categorised based on the prevalence of thin disc components, as we can see in the left panel of Fig. 3 where we can appreciate the stellar mass fraction of the cold disc at $z=0$ against that at $z=1$. The figure also shows a small evolution in the thick disc for the different samples, except for some lenticulars, which increased by a factor or two. When we compare the mass ratio between the thin disc and the thick disc, we see that this ratio increases with time for discs whereas, for lenticular and spheroidal, this ratio decreases, mainly because of the disruption of the cold disc. Finally, the right panel of Fig. 3 show the mass fraction of the classical bulge component that for spheroidal descendants, this increases with time reaching values close to $1$ whereas for disc galaxies, it increases up to $0.25$. Based on the analysis of the descendant samples, it is observed that $51$ per cent of disc galaxies at redshift $z=1$ maintain their cold disc component at redshift $z=0$, which accounts for $30$ per cent of the disc galaxies in the local universe if we take the number of discs at $z=0$ (349) of Rosas-Guevara et al. (2022) for the TNG50. Additionally, it is found that $26$ per cent of the $z=1$ disc galaxies undergo a transformation into spheroidal galaxies, while the remaining $23$ per cent transition into lenticular galaxies (see Table 1).
@@
It is also observed that a large fraction of disc galaxies ($49$ per cent) at redshift $z=1$ undergo a morphological transition. This transformation is likely attributed to external factors, such as tidal interactions or stripping resulting from mergers. Further details on this matter will be discussed in the subsequent section. Consequently, this phenomenon will have an impact on the pre-existing bar structures. In the instance of the disc descendants, it is observed that a subset of the discs had the ability to develop or dissolve their bar.
@@
Table 1: Descendant galaxy samples defined according to their final morphology ($z=0$). From left to right: Name sample, thin disc mass fraction $(D/T)_{\rm thin}$ at $z=1$ ($(D/T)_{\rm thin,z1}$), $(D/T)_{\rm thin}$ at $z=0$ ($(D/T)_{\rm thin,z0}$), number of galaxies, the fraction with respect to the parent disc sample at $z=1$ ($f_{\rm gal}$), number of bars at $z=1$ ($n_{\rm bar,z1}$), number of bars at $z=0$ ($n_{\rm bar,z0}$), number of bars formed after $z=1$ $n_{\rm bar,form,z<1}$, number of dissolved bars ($n_{\rm bar,lost,z<1}$), number of unbarred galaxies ($n_{\rm{unbar},z=1}$) and number of unbarred galaxies at $z=0$ ($n_{\rm{unbar},z=0}$).
@@
Figure 2: The $D/T$ distribution of disc galaxies at $z=1$ (top panel) and their descendants at $z=0$ (bottom panel). Note there was a morphological transformation overall in the population of the descendants of parent disc galaxies, especially, the thin disc component.
Figure 3: The evolution of the cold/thin disc, warm/thick disc and classical bulge components since $z=1$ for the three descendant subsamples: discs, lenticular and spheroidal descendants.

### 07 Feb 2024 v2

@@ 2.3 Identification of bars
Non-axisymmetric structures are identified by Fourier decomposing the face-on stellar surface density (e.g. Athanassoula & Misiriotis 2002; Zana et al. 2018a; Rosas-Guevara et al. 2020). We focus on $A_{2}(R)$, which is defined by the ratio between the second and zero terms of the Fourier expansion and its phase $\Phi(R)$ (see equations 1 and 2 in RG22). Both quantities $A_{2}(R)$ and $\Phi(R)$ have previously been used to characterise a bar structure, where the bar strength is defined as the value of the peak of $A_{2}(R)$, $A_{\rm 2,max}$. The phase should be constant inside the bar. We define a constant phase by calculating the standard deviation ($\sigma$) of $\Phi(R)$, including each time a new cylinder shell, and imposing $\sigma\leq 0.1$. We then define the bar extent ($r_{\rm bar}$) as the maximum radius where $\sigma\leq 0.1$, and the A2 profile first dips at $0.15$ or the minimum value of $A_{2}(R)$. As large values of $A_{2}(R)$ could also be due to transient events, such as mergers and interactions, we conservatively assume the bar is a long-lasting feature if the following three criteria are fulfilled.
@@
    The maximum of $A_{2}$ plus $A_{\rm 2,max}$ is greater than $0.2$.
@@
    The estimated age of the bar is greater than the time between the analysed output and the two previous simulation outputs ($0.33$ Gyr at $z=0$ and $0.17$ Gyr at $z=4$).
At lower redshifts, this latter criterion may exclude recently produced bars but remains consistent with the bar-selection criteria used by RG20 when the presence of a bar structure was observed in earlier outputs.
In addition to the criteria used to identify a bar structure, we need to determine the formation time of a bar, $t_{\rm bar}$. To this end, we retrace the evolution of $A_{\rm 2,max}$ up to the lookback time when $A_{\rm 2,max}\leq 0.2$ and $r_{\rm bar}(t_{\rm bar})\leq r_{\rm min}$ for more than two snapshots. In addition, the relative difference between the $A_{\rm 2,max}$ at a given snapshot and the $A_{\rm 2,max}$ at the two prior snapshots must not exceed $0.45$ within this time frame. This provides additional assurance that the bars we detect are stable structures.
@@
The catalogue of RG22 consists of 260 disc galaxies at $z=1,$ of which $125$ (0.48±0.03plus-or-minus0.480.030.48$\pm$ 0.030.48 ± 0.03444These are the binomials errors on the bar fractions, which are calculated using the number of galaxies and bars as $\sigma=(f_{\rm bars}(1-f_{\rm bars})/n_{\rm discs})^{0.5}$ ) have a bar. In this work, we ignore the galaxies that merged with a larger galaxy. This condition reduced our sample to $204$ disc galaxies at $z=1,$ including $93$ barred galaxies ($0.45\pm 0.03$). The details of the total sample can be found in Table 1. We follow the evolution of these galaxies from $z=1$ to $z=0$. Hereafter, we tag the $z=0$ sample as descendant galaxies.
@@ 2.4 Descendant galaxy samples
In order to ascertain the primary mechanisms influencing the disc galaxies at $z=1$, we employ the Sublink merger tree technique (Rodriguez-Gomez et al. 2015) to track their evolution over time. The parent disc sample was partitioned into three distinct subsamples based on the final morphology of their descendant galaxies at $z=0$. Our analysis specifically centres on the stellar mass fraction of the cold-disc component at $z=0$ ($(D/T)_{\rm thin,z0}$), as our initial dataset mostly consists of galaxies with a prominent cold-disc component. The division was made in the following manner:
  • Disc descendants refer to those galaxies at $z=1$ and their descendant galaxies at $z=0$ that have a massive cold-disc component, defined as $(D/T)_{\rm thin,z0}\geq 0.5$ and $(D/T)_{\rm thin,z1}\geq 0.5$.
  • Lenticular descendants are defined as those galaxies at $z=1$ that have a massive cold-disc component $(D/T)_{\rm thin,z1}\geq 0.5$. These galaxies have descendants at $z=0$ that possess an intermediate cold-disc component, specifically falling within the range of $0.50.3\geq(D/T)_{\rm thin,z0}<$ 0.5.
  • Spheroidal descendants defined as those galaxies at $z=1$ that have a massive cold-disc component $(D/T)_{\rm thin,z1}\geq 0.5$. However, their descendants at $z=0$ have seen a reduction in their cold-disc component, with values smaller than $(D/T)_{\rm thin,z0}=0.3$.
The disc descendants comprise $104$ galaxies representing $51$ per cent of the total parent sample, of which 36 galaxies are barred at $z=1$. Among these 104 galaxies, $47$ are classified as lenticular descendants ($23$ per cent of the parent samples) with $26$ barred galaxies at $z=1$. Finally, we get $53$ galaxies classified as spheroidal descendants, corresponding to $26$ per cent of the parent sample. The details can be seen in Table 1. In the evolution of the bar structures, we only focus on disc and lenticular galaxies.
@@ 2.5 Quantifying the effects of environment
The main aim of this study is to conduct a quantitative analysis of the impact of environmental factors on the development and progression of bars. Previous studies have provided evidence that mergers could facilitate the formation of a bar structure (e.g. Elmegreen, Elmegreen, & Bellin 1990; Łokas 2018). On the other hand, alternative studies have demonstrated that mergers can lead to the disruption of pre-existing bars (e.g. Zana et al. 2018b). Tidal interactions, such as those resulting from galaxy–galaxy interactions and flybys, are an additional external mechanism that can contribute to the formation of bars (e.g. Elmegreen, Elmegreen, & Bellin 1990; Lang, Holley-Bockelmann, & Sinha 2014; Łokas 2018; Peschken & Lokas 2019; Izquierdo-Villalba et al. 2022). The aim of the present work is to quantitatively assess three environmental factors —including the number of satellites as a function of time— that may potentially influence the evolution and production of bars. By identifying these external mechanisms, we seek to gain a deeper understanding of their impact on bars.
@@ 2.5.1 Merger histories and their influence
Major mergers are characterised by a stellar mass ratio of the secondary galaxy to the primary galaxy, denoted $\mu$, that surpasses $0.25$, whereas minor mergers are defined by $\mu$ values ranging from $0.01$ to $0.25$  [^5] The lower limit on the mass ratio for minor mergers ensures that both galaxies are properly followed. As our galaxy sample has stellar masses of $>10^{10}\rm M_{\sun}$, the secondary galaxy will have $M_{*}>10^{8}\rm M_{\sun}$ , which corresponds to above $1000$ times the initial mass of the gas cells.. In order to quantify the parameter $\mu$, the maximum stellar masses of the galaxies in preceding snapshots are used, namely those snapshots in which both galaxies are recognised as distinct structures by the SUBLINK algorithm.
According to McAlpine et al. (2020), the period where mergers can strongly affect a galaxy can be defined as the period during which galaxies are expected to undergo a close merger within two dynamical timescales of the halo relative to the desired lookback time (time of bar formation, time of bar dissolution). In this particular scenario, our focus lies on the temporal interval during which the bar undergoes formation or dissolution. The time period influenced by the merging process can be parameterised as
@@
where $t_{\rm obs}$ is the lookback time of formation or dissolution of the bar, $t_{\rm m}$ is the lookback time of the merger, and the dynamical time of the halo is defined as the free-fall time of a dark matter halo, that is,
$t_{\rm dyn,obs}\equiv\left(\frac{3\pi}{32G(200\rho_{\rm crit})}\right)^{1/2}.$ (2)
This dynamical time of the halo depends on redshift. As a point of reference, the dynamical time of a halo for three redshifts is approximately $0.9$ Gyr at $z=1$, $1.2$ Gyr at $z=0.5$, and $1.6$ Gyr at $z=0$. Positive values of $\eta_{\rm dyn,obs}$ indicate that the closest merger has already occurred, whereas negative values indicate that it will occur in the future. Absolute values of $|\eta_{\rm dyn,obs}|<1$ indicate that the merger takes place in one dynamical time. Therefore, we define the merger influence region for galaxies where the nearest merger occurs as $|\eta_{\rm dyn,obs}|<2,$ whereas values higher than $|\eta_{\rm dyn,obs}|>2$ are not influenced by a merger. Here, we focus mainly on major mergers, as they have the greatest potential for significant impact. However, we also calculated this for minor mergers.
@@ 2.5.2 Close massive companions
In a similar vein, to measure the likelihood that a galaxy will encounter a flyby event (not necessarily leading to a merger event), it is necessary to compute the proximity of the nearest large companion, where the ratio of their stellar masses exceeds $1/4$. Galaxies that may experience an influence due to a nearby encounter will possess nearby massive partners situated within a maximum distance of $d_{\rm cp}<100$ kpcs (Łokas 2021). Conversely, galaxies that are usually isolated are characterised by the closest massive companion being located at a distance of greater than $d_{\rm cp}>500$ kpc.
@@ 3 Morphological evolution of $z=1$ disc galaxies
Figure 1: Stellar mass of their descendants at $z=0$ as a function of the stellar mass of disc galaxies at $z=1$. Different markers and colours represent different morphologies of the descendant galaxies, as indicated in the legend. The black solid line represents a one-to-one relation. Panels along the margins show the distributions of stellar mass at $z=0$ and $z=1$. Dotted lines represent the median of each distribution and the black dotted line is the median stellar mass of the total sample. There is a mild evolution in stellar mass in the last $8$ Gyr in the parent disc sample.
@@
Before analysing the evolution of the bar structures, we focus on the morphological evolution of our parent sample, as defined in section 2.2. We investigated the evolution of our parent disc galaxies at $z=1$ by employing the tree merger algorithm outlined in section 2.1. The primary focus of our analysis is on the galaxy descendants of these disc galaxies at $z=0$.
@@
Figure 1 displays the scatter plot of the stellar mass of the parent disc sample at $z=1$ versus the stellar mass of their descendants at $z=0$. We can appreciate from the figure that, overall, there was an increase in stellar mass, although this growth is modest, with a median stellar mass of $10^{10.33}\rm M_{\sun}$ at $z=1$ rising to $10^{10.60}\rm M_{\sun}$ at $z=0$. We also note that some outliers, especially from spheroidal descendant galaxies, deviate from the overall distribution. The stellar mass of these galaxies decreases due to the entrance into a densely populated region, where they probably experience tidal stripping, as we show in the following section. This is confirmed when we look at the stellar-mass distributions (sided small panels) of the descendant galaxies split by shape, where spheroidal descendant galaxies have a smaller median stellar mass at $z=0$ than disc and lenticular descendant galaxies, whereas the parent sample at $z=1$ shows that the progenitors of the lenticular and spheroidal galaxies were more massive.
@@
It is important to note that the morphological evolution of our parent sample exhibits significant divergence. Figure 2 displays the distribution in the disc stellar mass fraction, denoted $D/T$ and represented in grey. The median $D/T$ decreases from $0.88$ at redshift $z=1$ to $0.75$ at redshift $z=0$. The dominant contribution is attributed to the thin-disc component, which exhibits a median of $(D/T)_{\rm thin,z1}=0.67$ at $z=1$. Nevertheless, when considering the case where $z=0$, it can be shown that the distribution of $(D/T)_{\rm thin,z0}$ in the cold-disc component is wider, spanning from $0.0$ to values higher than $0.75$. The median value of $(D/T)_{\rm thin,z0}$, is found to be $0.50$. This wider distribution of the cold component is also evident in the overall distribution of the discs, including all components. It should be noted that there is a lack of substantial evolutionary changes observed in the thick disc component, as indicated by a median value of $(D/T)_{\rm thin,z0,z1}\approx 0.2$ for both redshifts.
@@
In order to examine the changes in morphology within our parent sample, we have chosen three galaxy descendant samples as outlined in section 2.4 and Table 1. These samples have been categorised based on the prevalence of thin-disc components, as we can see in the left panel of Fig. 3 where we can appreciate the stellar mass fraction of the cold disc at $z=0$ against that at $z=1$. The figure also shows a small evolution in the thick disc for the different samples, except for some lenticulars that show an increase by a factor or two. When we compare the mass ratio between the thin disc and the thick disc, we see that this ratio increases with time for discs, whereas for lenticular and spheroidal galaxies, this ratio decreases, mainly because of the disruption of the cold disc. Finally, the right panel of Fig. 3 shows the mass fraction of the classical bulge component for spheroidal descendants; this increases with time, reaching values of close to $1,$ whereas for disc galaxies, it increases up to $0.25$. Based on the analysis of the descendant samples, it is observed that $51$ per cent of disc galaxies at redshift $z=1$ maintain their cold-disc component at a redshift of $z=0$, which accounts for $30$ per cent of the disc galaxies in the local Universe if we take the number of discs at $z=0$ (349) of Rosas-Guevara et al. (2022) for the TNG50. Additionally, we find that $26$ per cent of the $z=1$ disc galaxies transform into spheroidal galaxies, while the remaining $23$ per cent transition into lenticular galaxies (see Table 1).
@@
We also observe that a large fraction of disc galaxies ($49$ per cent) at redshift $z=1$ undergo a morphological transition. This transformation is likely attributed to external factors, such as tidal interactions or stripping resulting from mergers. Further details on this matter are discussed in the following section. Consequently, this phenomenon will have an impact on the pre-existing bar structures. Regarding disc descendants, we observe that a subset of the discs have the ability to develop or dissolve their bar.
@@
Table 1: Descendant galaxy samples defined according to their final morphology ($z=0$). From left to right: Sample name, thin-disc mass fraction $(D/T)_{\rm thin}$ at $z=1$ ($(D/T)_{\rm thin,z1}$), $(D/T)_{\rm thin}$ at $z=0$ ($(D/T)_{\rm thin,z0}$), number of galaxies, the fraction with respect to the parent disc sample at $z=1$ ($f_{\rm gal}$), number of bars at $z=1$ ($n_{\rm bar,z1}$), number of bars at $z=0$ ($n_{\rm bar,z0}$), number of bars formed after $z=1$ $n_{\rm bar,form,z<1}$, number of dissolved bars ($n_{\rm bar,lost,z<1}$), number of unbarred galaxies ($n_{\rm{unbar},z=1}$), and number of unbarred galaxies at $z=0$ ($n_{\rm{unbar},z=0}$).
@@
Figure 2: $D/T$ distribution of disc galaxies at $z=1$ (top panel) and their descendants at $z=0$ (bottom panel). We note that there was a morphological transformation overall in the population of the descendants of parent disc galaxies, especially for the thin disc component.
Figure 3: Evolution of the cold and thin disc, warm and thick disc, and classical bulge components from $z=1$ for the three descendant subsamples: disc galaxies, lenticular galaxies, and spheroidal descendants.